OpenStudy (anonymous):

what is the projection of (2,5) onto (1,3)? a. 1.7(2,5) b. 1.7(1,3) c. 0.5(1,3) Please help I don't know what it means by projection

OpenStudy (anonymous):

do you know the dot product of the vectors

OpenStudy (anonymous):

no

OpenStudy (anonymous):

for vectors a.b = |a||b|cos x where x is the angle between a and b

OpenStudy (anonymous):

okay so how do I use that to solve the problem?

OpenStudy (anonymous):

suppose we want projection of b on a then it is given by |b|cosx which equals (a.b)/|a|

OpenStudy (anonymous):

I'm sorry I still don't understand how this applies to my problem? Is (2,5) a and b or is (1,3)?

OpenStudy (anonymous):

is your question related to vectors

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then a is (2,5) and b is (1,3)

OpenStudy (anonymous):

a.b=2.1 +5.3 =17

OpenStudy (anonymous):

wait so are you saying (2*5)/ 2 equals either 1.7 or 0.5?

OpenStudy (anonymous):

but here we need (a.b) / |b| here |b| = sqrt(1^2 +3^2)= sqrt(10)

OpenStudy (anonymous):

okay i got 5.3

OpenStudy (anonymous):

the answer is 1.7(1,3)